what is (x+3)(x -2)??

Grade:9

9 Answers

Pankaj
9 years ago
Hello Student,

Please find the answer to your question

(x + 3)(x – 2) = x^2+x-6

Hope that you will be satisfied with this
Bhawna
9 years ago
ok. But can you show me the steps also.
Pankaj
9 years ago
x(x – 2) + 3(x – 2)
= x^2 – 2x + 3x – 6
= X^2 + x – 6

Hope that you will understand now

Thanks
Siddharth
9 years ago
Sir, what type of question is this? Can you explain this to me?
Bhawna
9 years ago
Hi Siddharth,
 
This is multiplication of linear factors
Pankaj
9 years ago
absolutely correct.

x(x – 2) + 3(x – 2)
= x^2 – 2x + 3x – 6
= X^2 + x – 6

Hope that you both have understood now
Siddharth
9 years ago
yes sir. Thank you
Bhawna
9 years ago
hello sir,
 
can you explain this to me one more time?
Pankaj
9 years ago
(x + 3)(x – 2)
=x(x – 2) + 3(x – 2)
= x^2 – 2x + 3x – 6
= X^2 + x – 6

Hope that you will understand now